OpenStudy (anonymous):

help pls! Find a ⋅ b. a = <2, 4>, b = <2, 5>

OpenStudy (smilewithmehun):

Are the ">" and "<" meant to be parenthesis?

OpenStudy (anonymous):

nope, they're for dot products

OpenStudy (anonymous):

of vectors

OpenStudy (jdoe0001):

\(\bf <a,b>\cdot <c,d>\implies a\cdot c+b\cdot d\)

OpenStudy (anonymous):

so is the answer <4,20> ?

OpenStudy (anonymous):

lol blaze it

OpenStudy (jdoe0001):

heh

OpenStudy (jdoe0001):

the answer for a dot product is just a number

OpenStudy (anonymous):

oh okay, so 24?

OpenStudy (anonymous):

thanks!

OpenStudy (jdoe0001):

yw